The position of an object moving along a straight line is given by x = 3 − 2t

2 + 4t

3 where x is

in meters and t in seconds.

a) Derive the expressions for the velocity and acceleration of the object as a function of time.

b) Find the position of the object at t = 0, t = 2s, t = 4s.

c) Find the displacement or the object between t = 2s and t = 4s; between t = 0s and t = 4s.

d) Find the average velocity between t = 2s and t = 4s; between t = 0s and t = 4s;

between t = 1s and t = 3s.

e) What is the instantaneous velocity at t = 2s? at t = 5s?

f) At what time(s) is/are the instantaneous velocities zero?

g) When does the instantaneous velocity have a maximum or a minimum value?

h) Find the change in velocity between t = 2s and t = 5s.

i) Find the average acceleration between t = 2s and t = 5s; between t = 1s and t = 3s.

j) Find the instantaneous acceleration of the object at t = 2s; t = 5s.


1
Expert's answer
2020-11-26T09:42:09-0500

Position is given by, "x=3-2t^2+4t^3"


Velocity is given by, "v=\frac{dx}{dt}=-4t+12t^2"


Acceleration is given by, "a=\frac{dv}{dt}=-4+24t"


Put value of t to find the position at any time,

Position at t=0, "x=3m"

Position at t=2, x=27m

Position at t=4, x=227m


Displacement is difference of final position and initial position.

"x(t=2-t=4)=227-27=200m"

"x(t=0-t=4)=227-3=224m"



Average velocity is total displacement per unit total time.

"v(t=4-t=2)=(200)\/2=100m\/s"

"v(t=0-t=4)=(224)\/4=55m\/s"

"v(t=3-t=1)=(93-5)\/2=44m\/s"


Velocity at any time is given by simply putting value of time.

"v(t=2)=40m\/s"

"v(t=5)=280m\/s"


Time when velocity is zero, "-4t+12t^2=0 \implies t=0, 1/3s"


For velocity to be maximum or minimum, first derivative of the velocity must be zero.

"-4+24t=0 \implies t=6s"


Change in velocity in t=2 - t=5 s, "v=280-40=240m\/s"


Average acceleration is total change in velocity per unit total time.

"a(t=2-t=5)=(240)\/3=80m\/s^2"

"a(t=1-t=3)=(96-8)\/(3-1)=44m\/s^2"


Instantaneous acceleration is given by simply putting value of time.

"a(t=2)=48-4=44m\/s^2"

"a(t=5)=120-4=116m\/s^2"